# Gentle-CLIP: Exploring Aligned Semantic In Low-Quality Multimodal Data With Soft Alignment

### v1

@@ 3.3 Objective loss
where $\mu$ is a hyperparameter and $\mathcal{L}_{GC}$ is used to guide the training process with accurate supervised alignment information rather than semantic distribution similarity, which is necessary for avoiding negative optimization. Meanwhile, if the data from different modalities can achieve augmentation according to the common semantics rather than the pattern in the single modality, the performance of related method may realize further growth[83].
The Overall Pretraining Objective: Our method aims to adopt matched pairs as well as unsupervised data in a batch at the same time. In this way, during the pretraining process, we can utilize comprehensive unsupervised data as well as the alignment constraint from matched pairs to realize robust and stable optimization process. Moreover, through samantic distribution alignment, the knowledge learned from unsupervised data and matched pairs can potentially interact with each other which could guide to enlarge the range of knowledge. For overall pretraining objective, we seek to minimize the loss functions of all pretraining tasks simultaneously:
@@
where $\theta$ denotes all learnable parameters in encoder networks and projection head networks. $\alpha$, $\delta$ and $\eta$ are hyperparameters which is used to control the influence of different pretraining tasks.
4 Experiments
In order to evaluate the effectiveness of our proposed method, we conduct extensive experiments in various specialized fields, including protein representation, remote sensing as well as general vision-language field. In addition, we analyze the roles of key modules by ablation experiments.
4.1 Quantitative Analysis About Sampling Size
As mentioned above, there exists implicit alignment information between different modalities with similar semantic distribution even if there is no definite matched pairs. Therefore, if we can acquire unimodal batches which reflect the real distribution of original data by stochastic sampling in each modality, it is derivable that each batch from different modalities also keeps similar semantic distribution and can be used for subsequent training process. Obviously, sampling size significantly influence the ability whether batches are on behalf of original distributions. Hence We attempt to quantitatively analyze the ability of different scales of sampling size for representing the original distribution, which can guide to choose the proper size. By applying soft Parzen-window method which is used for non-parametric density estimation, we can calculate the representing confidence of specific sample batch with given size. Through experimental verification, it could be concluded that sample batches will be able to represent original complicated distribution effectively when sampling size is over $64$. Furthermore, if different modal data is from the same semantic distribution, the batches with sampling size over $64$ will also keep the similar semantic distribution. More Detailed process of method and relevant analysis will be displayed in Appendix E.
Table 1: Benchmark results on protein representation field. Bold denotes the best results while underline represents the second best value. Two-stream networks outperform in most tasks and Gentle-CLIP can effectively explore latent alignment from unmatched data.
Input Method Gene Ontology Enzyme Commission Average
@@ 4.1 Quantitative Analysis About Sampling Size
4.2 Evaluation On Single Protein Function Prediction
Overview of tasks and training setup: To examine the efficacy of Gentle-CLIP in non-vision-language multimodal domains with insufficient alignment data, we conduct experiments in the protein representation field. Proteins can be defined using a multi-level structure and most previous works take aligned sequence and structure as input for single-stream network to capture the invariance features[26]. Due to limited aligned data, these intricately designed models struggle into trouble. Following [30], we consider sequence and structure as two modalities and apply Gentle-CLIP to realize multimodal fushion by pulling semantic distributions closer from extensive unsupervised data. Structure encoder is designed based on CDConv[23] while ESM-2[91] is selected as sequence encoder. We adopt CATH $4.2$ dataset for pretraining and this process lasts $100$ epochs. According to the same settings in [23], we evaluate the proposed method on the following four tasks: protein fold classification[26], enzyme reaction classification[26], gene ontology (GO) term prediction[27] and enzyme commission (EC) number prediction[27]. More details is shown in Appendix F and D.
Results: The performance of downstream tasks are shown in Table 1 while results of previous baselines are from [23] and [1]. Thereinto, CLIP(1/2 CATH) is a two-stream network with 50% CATH data for pretraining while CLIP(CATH) is pretrained on the whole CATH datasets. Gentle-CLIP(Ours) adopts 50% CATH data as supervised pairs while the rest are considered as unlabeled data. Moreover, we add an Average item to evaluate the overall performance. We first verify the effect of two-stream network compared to single-stream model and the results are displayed at the last line in Table 1. We can find that CLIP achieve better results at most downstream tasks and show superiority especially in EC number prediction. Further, it is obvious that Gentle-CLIP dramatically narrow the overall gap between CLIP pretrained on the whole CATH and the performance is even better at a few downstream tasks. This may be due to the fact that Gentle-CLIP can explore implicit alignment through the fine-grained semantic distribution constraint of SDD while CLIP only focuses on local representation match which may loss some global distribution information.
4.3 Evaluation On Remote Sensing Datasets
Overview of tasks and training setup: The models in remote sensing field can acquire comprehensive knowledge by jointly learning satellite images and corresponding captions. However, the training datasets are usually composed of web-crawled data without integrality and annotating captions may also need various expert knowledge, which can be expensive and time-consuming. So it is essential to evaluate the performance of Gentle-CLIP on limited matched pairs which is hard to tackle by traditional methods. Following [28], Gentle-CLIP is pretrained on the union of RSICD[55], UCM[56] and Sydney[57] with zero-shot classification and image-text retrieval as downstream tasks. ResNet[2] and transformer[5] are chosen as single modal encoder and Gentle-CLIP is pretrained for $25$ epochs. We subsample $10\%$ of image-text pairs for supervised learning while the remaining data is served unlabeled but conform to the same semantic distribution. Similarly, Top-1 classification accuracy is used to evaluate the performance on zero-shot classification while recall is applied for image-text retrieval tasks. Detailed descriptions will be presented in the Appendix C and D.
Table 2: Benchmark results on remote sensing field. Bold is the best average results and Gentle-CLIP improves the performance at most datasets through learning from unsupervised data.
@@ 4.3 Evaluation On Remote Sensing Datasets
Gentle-CLIP(ours) 69.2$\pm$0.8 67.5$\pm$1.1 89.0$\pm$1.6 66.2$\pm$0.9 76.2$\pm$0.9
Results: Table 2 displays the results of zero-shot classification and the first five items are from [28]. In this experiment, Gentle-CLIP is designed based on S-CLIP and trained to narrow the embedding distribution gap between paired texts and unlabeled images under the guidance of SDD. The whole distributions of batches from different modalities may keep similar even though the explicit matching relationship is unknown between specific samples. For zero-shot classification, our method shows outstanding performance except RSSCN7. We can also find that existing methods are all difficult to bring much gain compared with other datasets, so it is believed that RSSCN7 may have significant gap with training set resulting in greater difficulty for inference. As shown in Appendix G, Gentle-CLIP consistently improves the results in image-text retrieval, which proves that less distribution gap between texts and unlabeled images brings robust pseudo-labels and stable distribution structure.
4.4 Evaluation On General Vision-Language Retrieval
Overview of tasks and training setup: Besides above specialized domains, we also evaluate the performance of Gentle-CLIP in general vision-language field. Experiments are carried on Flickr-8k[20] and Mini COCO while image-text retrieval is adopted as downstream task. Moreover, the vision encoder employs ResNet-50 or ViT-32[24] and BERT[25] acts as the text encoder. We choose the first description of each image as corresponding caption while dropout ratio is set to $0.3$ for augmentation following [40]. Pretraining process lasts $50$ epochs and batch size is $64$ with learning rate equalling to $0.001$. Furthermore, $1/3$ multimodal data is considered to be matched while the rest data is unlabeled. Datailed statements of datasets will be displayed in Appendix C.
Table 3: Benchmark results on general vision-language field. CLIP(1/3) is the baseline while values highlighted in green indicate the improvements. Gentle-CLIP enlarges the knowledge range and especially brings advantages in Mini COCO dataset with VIT as image encoder.
@@ 4.4 Evaluation On General Vision-Language Retrieval
Gentle-CLIP +2.1 +2.9 +1.9 +3.6 +4.3 +12.3 +5.1 +12.6
@@
Results: We evaluate the performance with different models as well as datasets and the results of image-text retrieval is shown in Table 3. CLIP(1/3) is trained by 1/3 dataset while CLIP(1) learns alignment on the whole dataset. Gentle-CLIP employs 1/3 matched data for supervised learning and enlarge knowledge range from rest unpaired data. For better observing effects of different strategies, we adopt the improvement value as yardstick and baseline is CLIP(1/3). We can find that Gentle-CLIP continuously brings gains regardless of settings and the overall performance of VIT is better than ResNet for given datasets. From the experimental results, we may conclude that multimodal fushion could be simple when different modality models have similar distance measurement.
4.5 Ablation
Gentle-CLIP applies different kinds of objectives to explore implicit semantic alignment from low-quality multimodal data and shows excellent performances in various experiments. Then, we will further analyze the internal mechanism by replying to the following noteworthy questions.
Q1: How will SDD and SSL influence the final effectiveness? To answer this question, we conduct experiments in protein representation field and evaluate the performance at protein fold classification as well as enzyme commission number prediction. We focus on the effects of these two objectives and results are shown in Table 4.5. We can find that Gentle-CLIP trained with both objectives shows better performance except superfamily classification while the model only trained by SDD outperforms at this item but fall into negative optimization in EC number prediction. Single SSL brings weak improvement compared with the baseline so it also acquires additional alignment information during pretraining process. Through above results, we believe that the combination of these two objectives brings more advantages rather than simple stack of respective effect, in other words, these two losses interact and depend on each other. To be specific, SDD can exploit alignment information in unsupervised data but may cause mode collapse and negative learning as shown at the third line. While adding SSL can further constrain the optimization direction and increase the stability of the overall distribution. However, SSL will also affect the latent distribution of similar samples, hence it is necessary to balance the relationship to achieve better performance[83].
Q2: How is the performance if we change some modules of SDD? With fine-grained distribution similarity measurement, SDD plays a critical role in semi-supervised fushion. So it is essential to deconstruct SDD and analyse which settings may lead to better effects. We make retrieval experiment on Flickr-8k with Top-3 recall and Table 4.5 display the results. We focus on how to calculate the distance between samples as well as the difference between distributions. Relative distance(RD) $\|x-t_{i}\|^{2}/\sigma(T)$ in eq. 5 can eliminate the indistinguishability in tight latent cluster compared to absolute distance while Kullback-Leibler Divergence(KL) in eq. 4 may be more suitable for distribution contrast with MSE. It is clear that models achieve the greatest improvement when adding RD and KL simultaneously. Moreover, if improved Softmax function in [47] can be employed to modify the distance $\|x-t_{i}\|^{2}$ with different attention, we will control distribution alignment granularity through different temperatures. Other ablation results will be shown in Appendix I.
@@ 4.5 Ablation
5 Conclusion
We propose a semi-supervised multimodal alignment method named as Gentle-CLIP and design a novel objective SDD which can measure the fine-grained difference between latent distributions. Besides, we introduce other loss functions to realize rubost generalization through the analysis of latent embedding. We demonstrate the superiority of our method across various fields and provide a possible way for multimodal fushion in specialized domains with insufficient aligned data by exploring latent alignment information from unimodality.
Appendix A The analysis of latent embedding
We hold the opinion that the ultimate embedding is a combination representation of modality, structure as well as semantic. If data is from different modalities, the modality representation may have difference and lead to the gap between ultimate embeddings. Meanwhile, if the data is from single modality, the representation may be separated due to the gap of semantic rather than modality. In addition, even though the data keeps similar semantic and are from the same modality, the embedding may also have difference because of the internal structural gap. The visualization of relationship between these three representations are shown in Figure 3.
Figure 3: The visualization of relationship between different representations.
Appendix B Difference between SDD and Sinkhorn Algorithm
For Sinkhorn Algorithm, we will analyse its role on the basis of SuperGlue[84]. In the attention-based graph neural network processing, each pair of keypoints (one from image A and one from image B) gets a combined feature representation. SuperGlue constructs a score matrix $S$ by calculating the inner product of these feature representations, where each element $S_{i,j}$ indicates the matching score between the $i$-th keypoint in image A and the $j$-th keypoint in image B. In order to handle the unmatched keypoints elegantly, it introduces an additional "dustbin" to the score matrix, allowing keypoints which find no match from the other image to be assigned to this dustbin.
The score matrix $S$ is transformed using the exponential function to obtain $K=e^{S}$. Sinkhorn algorithm iteratively adjusts $K$ through alternating row and column normalizations:

### v2

@@ Introduction
where $\mu$ is a hyperparameter and $\mathcal{L}_{GC}$ is used to guide the training process with accurate supervised alignment information rather than semantic distribution similarity, which is necessary for avoiding negative optimization. Meanwhile, if the data from different modalities can achieve augmentation according to the common semantics rather than the pattern in the single modality, the performance of related method may realize further growth(Huh et al. 2024).
The Overall Pretraining Objective: Our method aims to adopt matched pairs as well as unsupervised data in a batch at the same time. In this way, during the pretraining process, we can utilize comprehensive unsupervised data as well as the alignment constraint from matched pairs to realize robust and stable optimization process. Moreover, through semantic distribution alignment, the knowledge learned from unsupervised data and matched pairs can potentially interact with each other which could enlarge the range of knowledge. For overall pretraining objective, we seek to minimize the loss functions of all pretraining tasks simultaneously:
@@
where $\theta$ denotes all learnable parameters in encoder and projection head networks. $\alpha$, $\delta$ and $\eta$ are hyperparameters used to control the impacts of different pretraining tasks.
Experiments
In order to evaluate the effectiveness of proposed method, we conduct extensive experiments in various fields, including protein representation, remote sensing as well as general vision-language field. In addition, we design sufficient ablation experiments to analyze the roles of key modules.
Quantitative Analysis About Sampling Size
As mentioned above, there exists implicit alignment information between different modalities with similar semantic distribution even if there is no definite matched pairs. Therefore, if we can acquire unimodal batches which reflect the real distribution of original data by stochastic sampling in each modality, it is derivable that each batch from different modalities also keeps similar semantic distribution and can be used for subsequent training process. Obviously, sampling size significantly influence the ability whether batches are on behalf of original distributions. Hence We attempt to quantitatively analyze the ability of different scales of sampling size for representing the original distribution, which can guide to choose the proper size. By applying soft Parzen-window method, we can calculate the representing confidence of sample batch with given size. Through experimental verification, it could be concluded that sample batches will be able to represent original complicated distribution effectively when sampling size is over $64$. Furthermore, if different modal data is from the same semantic distribution, the batches with sampling size over $64$ will also keep the similar semantic distribution. More Detailed process of method and relevant analysis will be displayed in Appendix E.
Method Gene Ontology EC Average
@@
Table 1: Benchmark results on protein representation field. Bold denotes the best results while underline indicates the second best value. Two-stream networks like CLIP(CATH) outperform in most tasks and Set-CLIP can effectively reduce the gap between CLIP(CATH) with only half paired data of CATH for supervised alignment.
@@
Set-CLIP(ours) 69.2$\pm$0.8 67.5$\pm$1.1 89.0$\pm$1.6 66.2$\pm$0.9 76.2$\pm$0.9
Table 2: Benchmark results on remote sensing field. Bold is the best average results and Set-CLIP improves the performance at most datasets through learning from unsupervised multimodal data.
Evaluation On Single Protein Function Prediction
Overview of tasks and training setup: To examine the efficacy of Set-CLIP in non-vision-language multimodal domains with insufficient alignment data, we conduct experiments in the protein representation field. Proteins can be defined using a multi-level structure and most previous works take aligned sequence and structure as input for single-stream network to capture the invariance features(Hermosilla et al. 2020). Due to limited aligned data, these intricately designed models struggle into trouble. Following (Zheng et al. 2023a), we consider sequence and structure as two modalities and apply Set-CLIP to realize multimodal fusion by pulling semantic distributions closer from extensive unsupervised data. Structure encoder is designed based on CDConv(Fan et al. 2022) while ESM-2(Lin et al. 2022) is selected as sequence encoder. We adopt CATH $4.2$ dataset for pretraining and this process lasts $100$ epochs. According to the same settings in (Fan et al. 2022), we evaluate the proposed method on the following four tasks: protein fold classification, enzyme reaction classification, gene ontology (GO) term prediction and enzyme commission (EC) number prediction(Gligorijević et al. 2021). More details of this experiment is shown in Appendix F and D.
Results: The performance of downstream tasks are shown in Table 1 while results of previous approaches are from (Fan et al. 2022; Zhang et al. 2023; Xu et al. 2023a). Thereinto, CLIP(1/2 CATH) is a two-stream network with 50% CATH data for pretraining while CLIP(CATH) is pretrained on the whole CATH datasets. Set-CLIP(Ours) adopts 50% CATH data as supervised pairs while the rest are considered as unlabeled data. Moreover, we add an Average item to evaluate the overall performance. We first verify the effect of two-stream network compared to single-stream model and the results are displayed at the last line in Table 1. We can find that CLIP achieve better results at most downstream tasks and show superiority especially in EC number prediction. Further, it is obvious that Set-CLIP dramatically narrow the overall gap between CLIP(CATH) and the performance is even better at a few downstream tasks. This may be due to the fact that Set-CLIP can explore implicit alignment through the fine-grained semantic distribution constraint of SDD while CLIP only focuses on local representation match which may loss some global distribution information.
Evaluation On Remote Sensing Datasets
Overview of tasks and training setup: The models in remote sensing field can acquire comprehensive knowledge by jointly learning satellite images and corresponding captions. However, the training datasets are usually composed of web-crawled data and annotating captions may also need various expert knowledge, which can be expensive and time-consuming. So it is essential to evaluate the performance of Set-CLIP on limited matched pairs which is hard to tackle by traditional methods. Following (Mo et al. 2023), Set-CLIP is pretrained on the union of RSICD, UCM and Sydney with zero-shot classification and image-text retrieval as downstream tasks. ResNet(He et al. 2016) and transformer(Vaswani et al. 2017) are chosen as encoders and Set-CLIP is pretrained for $25$ epochs. We subsample $10\%$ of image-text pairs for supervised learning while the remaining data is served unlabeled but conform to the same semantic distribution. Similarly, Top-1 classification accuracy is used to evaluate the performance on zero-shot classification while recall is applied for image-text retrieval tasks. More Details will be presented in the Appendix C and D.
@@
Set-CLIP +2.1 +2.9 +1.9 +3.6 +4.3 +12.3 +5.1 +12.6
Table 3: Benchmark results on general vision-language field. CLIP(1/3) is the baseline and values highlighted in green indicate the improvements. Set-CLIP brings benefits across diverse settings, especially in Mini COCO dataset with VIT as encoder.
Figure 3: Retrieval results of CLIP and Set-CLIP under completely unsupervised scenario.
Results: Table 2 displays the results of zero-shot classification and the first five lines are from (Mo et al. 2023). In this experiment, Set-CLIP is designed based on S-CLIP and trained to narrow the embedding distribution gap between unlabeled images and texts under the guidance of SDD. The whole distributions of batches from different modalities may keep similar even though the explicit matching relationship is unknown between specific samples. For zero-shot classification, our method shows outstanding performance except RSSCN7. We can also find that existing methods are all difficult to bring much gain compared with other datasets, so it is believed that RSSCN7 may have significant gap with training set resulting in greater difficulty for inference. As shown in Appendix G, Set-CLIP consistently improves the results in image-text retrieval, which proves that less distribution gap between unlabeled images and texts brings robust pseudo-labels and stable distribution structure.
Evaluation On General Vision-Language Retrieval
Overview of tasks and training setup: Besides above specialized domains, we also evaluate the performance of Set-CLIP in general vision-language field. Experiments are carried on Flickr-8k(Hodosh, Young, and Hockenmaier 2013) and Mini COCO while image-text retrieval is adopted as downstream task. Moreover, the vision encoder employs ResNet-50 or ViT-32(Dosovitskiy et al. 2020) and BERT(Devlin et al. 2018) acts as the text encoder. We choose the first description of each image as corresponding caption while dropout ratio is set to $0.3$ for augmentation following (Gao, Yao, and Chen 2021). Pretraining process lasts $50$ epochs and batch size is $64$ with learning rate equalling to $0.001$. Furthermore, $1/3$ multimodal data is considered to be matched while the rest data is unlabeled. Datailed statements of datasets will be displayed in Appendix C.
@@
Results: We evaluate the performances with different models as well as datasets and the results is shown in Table 3. CLIP(1/3) is trained by 1/3 dataset while CLIP(1) learns alignment on the whole dataset. Set-CLIP employs 1/3 matched data for supervised learning and enlarge knowledge range from rest unpaired data. For better observing effects of different strategies, we adopt the improvement value as yardstick and baseline is CLIP(1/3). We can find that Set-CLIP continuously brings gains regardless of settings and the overall performance of VIT is better than ResNet for given datasets. From the experimental results, we may also conclude that multimodal fusion could be simple when different modality models have similar distance measurement.
SSL SDD Fold Classification EC
Fold Super Family
@@
2613
58.5 80.1 99.6 0.878
59.1 79.7 99.6 0.884
Table 4: Ablation results evaluated on protein representation field to analyze the roles of SSL and SDD. Thereinto, Super denotes Superfamily task and EC is EC number prediction. The values in bold is the best result at each task.
Ablation
Set-CLIP applies different kinds of objectives to explore implicit semantic alignment from low-aligned multimodal data and shows excellent performances in various experiments. Then, we will further analyze the internal mechanism by replying to the following noteworthy questions.
Q1: How will SDD and SSL influence the final effectiveness? To answer this question, we conduct experiments in protein representation field and results are shown in Table 4. We can find that Set-CLIP trained with both objectives shows better performance except superfamily classification while the model only trained by SDD outperforms at this item but fall into negative optimization in EC number prediction. Single SSL brings weak improvement compared with the baseline so it also acquires additional alignment information during pretraining process. Through above results, we believe that the combination of these two objectives brings more advantages rather than simple stack of respective effect, in other words, these two losses interact and depend on each other. Specifically, SDD can exploit alignment information in unsupervised data but may cause mode collapse(Du et al. 2023) and negative learning as shown at the third line. While adding SSL can further constrain the optimization direction and increase the stability of overall distribution. However, SSL will also affect the latent distribution of similar samples, hence it is necessary to balance the relationship to achieve better performance(Huh et al. 2024).